Castro’s daily production function for producing custom-printed mugs is q = 15K^0.5*L^0.5. The price of labor is $40 and the price of capital is $10.

a. Find and sketch the expansion path.

b. How much capital and labor should Castro employ to produce 150 mugs a day?

c. What is the lowest possible cost of producing 150 mugs a day?

d. Repeat parts (a) and (b), for an output level of 450 mugs.

e. What is the firm’s LRAC at q = 150? At q = 450? What does the firm’s LRAC look like over this range of output? Are there economies of scale? Explain.
Answer #1

(a) q= 15k 0.5 10.5 09 = MP2 = 1510.5)ko .5 20.5-1 O2 MPL = 7.56 0.5,-0.5 Ok - MPx = 1510967 * 0.5-4,0.5 mPx = 7.5*0.5, 0.5 Mb) eh put - 9 = K=4L in proan 15kous 20.5 - 14.12.0-5 10-5 a = a 2005 Lois - a= 22 = 20 K= 4L = 4/2) = 4a 2=150 L= 150=5 L=5e) we have ha a x = 42 с 4) + К = yoq + 14422 - 402 4oz 802 30 30 LRATC = C = 84 | 1) LRATC - cet q = 150 LRATC= 8 LRATC= 700

Thus,there is constant economies of scale.